The Missou Manufacturing Company recorded overhead costs of $14,352 at an activity level of 5,050 machine hours and $7,902 at 2,470 machine hours. The records also indicated that overhead of $9,815 was incurred at 2,770 machine hours. What is the total estimated cost for 2,770 machine hours using the high-low method to estimate the cost equation?

$6,925.

$9,815.

$8,652.

$7,902.
Answer #1

Variable cost per unit=[Total cost at highest level-Total cost at lowest level]/(Highest level-Lowest level)

=(14352-7902)/(5050-2470)

=$2.5 per machine hour

Hence total fixed cost=14352-(2.5*5050)=1727

Hence total estimated cost=Total fixed cost+Total variable cost

=1727+(2.5*2770)

=$8652

NOTE:Total fixed costs and variable cost per unit do not change with change in units